# Music Foundation Model as Generic Booster for Music Downstream Tasks

### v2

@@ 4.2.3 Music Mixing: Mix-Wave-U-Net, CRAFx2
Mix-Wave-U-Net (Steinmetz et al., 2022) along with a modified CRAFx (Martínez-Ramírez et al., 2020), henceforth referred to as CRAFx2, are used as the baselines. The input to both networks is the stereo stems pre-processed by Fx-normalization (Martínez-Ramírez et al., 2022), and the output is the stereo mixture.
These models do not handle high-level information that is relevant to mixing, such as genre, instrumentation, or mood. We thus expect that conditioning these models with the extracted features, which implicitly contain such information, will improve mixing performance. The features are computed from the monaural downmix of the mixture, which corresponds to the summation of the Fx-normalized input stems. To incorporate these features, we condition both networks using Feature-wise Linear Modulation (FiLM) layers (Perez et al., 2018). For Mix-Wave-U-Net, we inject features into the up-sampling and bottleneck one-dimensional (1D) convolutional blocks. For CRAFx2, we use FiLM layers to condition both the latent-space mixer and synthesis back-end (see Appendix F.1).
We train all models on MUSDB18, and the stereo-invariant loss, along with all training hyperparameters, remains the same, as in a previous study (Martínez-Ramírez et al., 2022). Due to the inherent subjectivity of the task, identifying a single best model is challenging. Thus, as shown in Table 1, an objective evaluation is conducted by measuring the proximity between the output mixes and target mixes on the same test sets as in the above study (Martínez-Ramírez et al., 2022). The proximity measurement is based on objective metrics (Steinmetz et al., 2022; Martínez-Ramírez et al., 2022). These metrics consist of spectral, panning, dynamic, and loudness low-level audio features, which are related to the key audio characteristics often manipulated during the mixing process. Further details and experiments are provided in Appendices F.2 and F.3, respectively.
@@
Using the SoniDo features leads to the best model performance. The MusicGen features provide improvements but not as effective as those of SoniDo, particularly for CRAFx2, where the default model outperforms MusicGen in various metrics. Thus, we assume this performance gap of MusicGen may be attributed to its training on 32 kHz audio compared with the 44.1 kHz used for SoniDo, which could limit its effectiveness for full-band tasks, such as music mixing, that require higher sampling rates.
There is no data-driven approach that used task-agnostic features of the input stems for music mixing improvement. Thus, we can conclude that incorporating the SoniDo features benefits both the training and performance of automatic music mixing models. This aligns with recent design studies (Lefford et al., 2021; Vanka et al., 2023), advocating for the incorporation of contextual inputs.
Table 6: Music mixing. Evaluation results on MDXDB21-dry test set and MUSDB18 test set. Objective metrics include mean absolute percentage error for audio effect-related features, their average, and stereo-invariant loss. More detail provided in Table 15.
@@
Mix-Wave-U-Net (default) 0.234 0.215 0.073 0.168 0.173 89.63 0.201 0.164 0.085 0.167 0.154 34.25
@@
Mix-Wave-U-Net + SoniDo 0.226 0.180 0.067 0.131 0.151 78.18 0.186 0.175 0.063 0.179 0.151 30.11
CRAFx2 (default) 0.193 0.179 0.070 0.152 0.148 82.09 0.193 0.154 0.081 0.165 0.148 32.86
@@ 6 Conclusions
We extended the use of music foundation models from MIR to generic music downstream tasks. The task-agnostic intermediate representation extracted using our music foundation model SoniDo has been applied to task-specific models of music tagging, music transcription, music source separation, and music mixing. On the basis of the evaluation results, performance improvement is observed for all selected music downstream tasks. This suggests that incorporating the intermediate features extracted from a pre-trained music foundation model should be considered as a generic booster in future development of task-specific models. This is especially helpful when it is difficult to acquire a sufficient dataset or the computation resource does not allow large-scale training. A study on the bias propagation of a pre-trained music foundation model to a downstream task model should be conducted in another future work.
@@ Appendix B Architecture of SoniDo
SoniDo is a generative model that generates music samples conditioned on given text prompts. We train the model using an internal dataset consisting of $4,000$h of studio-quality library music sampled at $44.1$ kHz. To generate music samples, we use the three priors in SoniDo consecutively from the top level to the bottom level. It is important to note that during music generation, we use the token sequences sampled from top and middle priors for conditioning. For training and feature extraction, however, we rely on ground-truth tokens.
@@ B.3 Evaluation on Music Generation
For evaluation, we used the MusicCaps dataset (Agostinelli et al., 2023), which includes 5,521 ***We used a subset of 5,514 samples due to the presence of expired or invalid URLs in MusicCaps items. Efforts were made to retrieve the missing files; however, three samples are absent despite our efforts. pairs of audio clips and their text captions written by musicians. Subsequently, we used SoniDo to generate $10$-s audio samples, conditioned upon textual captions of MusicCaps. Following the objective evaluation methodology(Copet et al., 2023), we evaluated SoniDo’s performance on MusicCaps using three objective metrics: the Fréchet audio distance (FAD) (Kilgour et al., 2019), Kullback-Leibler Divergence (KL), and CLAP (Wu* et al., 2023) score.
@@
We used FAD to measure the overall quality of the generated samples. It is a reference-free metric that compares embedding statistics computed on an evaluation set with embedding statistics computed on the clean dataset. A lower FAD indicates a higher degree of similarity between the embedding statistics from generated audio samples and the embedding statistics from the ground-truth (i.e., MusicCaps). We used the official VGGish based-FAD computation module***https://github.com/google-research/google-research/tree/master/frechet_audio_distance to extract embeddings from audio. Since VGGish was trained using $16$-kHz audio, all the audio data were resampled to $16$ kHz.
@@
MusicGen-large (public ***Publicly released MusicGen-large trained on non-vocal dataset: https://github.com/facebookresearch/audiocraft/blob/main/model_cards/MUSICGEN_MODEL_CARD.md, ) 32kHz 5.88 1.41 0.27
@@ D.1 Piano Transcription with Linear Layers
The mel spectrogram is a 3D tensor ($B$, $N$, $F$), where $B=8$ is the batch size, $N=256$ is the number of frames, and $F=256$ is the number of frequency bins. A linear layer converts the tensor into ($B$, $N$, $Z$) along the last axis, where $Z=512$ is a common size for latent embeddings. The SoniDo and MusicGen features are formed as a 3D tensor ($B$, $V$, $G$), where $V$ is the number of tokens per $N$ frames (2,048 for the top layer of SoniDo, 8,192 for the middle layer of SoniDo, and 299 for MusicGen), and $G$ is the embedding size of features (4,800 for the top layer of SoniDo, 3,200 for the middle layer of SoniDo, 2,048 for MusicGen Large, and 1,024 for MusicGen Small). The first linear layer converts $G$ to $Z$, then the second linear layer converts $V$ to $N$ to make the tensor into the shape of ($B$, $N$, $Z$). The tensors are then concatenated along the last axis. If the top layer of SoniDo, middle layer of SoniDo, or MusicGen is used, the dimension of the last axis is $2Z$, whereas if both layers of SoniDo are used, the dimension is $3Z$. The following linear layer is used to convert $2Z$ or $3Z$ to $2P$, where $P=88$ is the number of pitches. Finally, the tensor is split into two ($B$, $N$, $P$) tensors followed by a sigmoid activation to indicate the estimated frame and onset information. We followed Kong et al. (2021) to extract the precise timing of onsets from datasets, in which we set the hyper-parameter to control the target sharpness $J$ to $3$. The loss function for frame and onset is binary cross entropy.
@@
(a) Linear transcription model using SoniDo
(b) hFT-Transformer using SoniDo
@@
We trained multiple models with the following input: (1) spectrogram only, (2) spectrogram and the top layer of the SoniDo features, (3) spectrogram and the middle layer of the SoniDo features, (4) spectrogram and both the top and middle layers of the SoniDo features, (5) spectrogram and the MusicGen Large features, and (6) spectrogram and the MusicGen Small features. For each model, we trained for 50 epochs on one A100 graphics processing unit (GPU), using Adam (Kingma & Ba, 2015) optimizer with a learning rate of 1e-4. PyTorch ReduceLROnPlateu was used for learning-rate scheduling with default parameters. We chose to use the checkpoint with the highest F1 score in the validation split for each model.
As listed in Table 3, the model using the SoniDo features or MusicGen features outperformed the spectrogram-only baseline. In particular, the contribution of SoniDo’s middle layer was higher than that of the top layer. These results suggest that the SoniDo features are promising for music transcription tasks.
@@ D.2 Linear Instrument-agnostic Music Transcription
The detailed model architecture of hFT-Transformer has been described in (Toyama et al., 2023). Figures 6(b) and (d) show the feature-spectrogram alignment scheme for hFT-Transformer. The feature tokens and spectrogram cannot be perfectly aligned due to the different sampling frequencies in hFT-Transformer ($16$ kHz), SoniDo ($44.1$ kHz), and MusicGen ($32$ kHz). The $N$ frame in the spectrogram corresponds to roughly $707$ tokens of the top layer of the SoniDo features, $,82$ 828 tokens of the middle layer of the SoniDo features, and $104$ tokens of the MusicGen features. Figure 7 shows the modified hFT-Transformer that accepts the SoniDo or MusicGen features as the additional input. Following the setup in D.1, the SoniDo and MusicGen feature tokens have three dimensions ($B$, $V$, $G$), where $B=8$ is the batch size, $V$ is the number of tokens per number of frames $N=128$, and $G$ is the embedding size of the feature tokens. The first linear layer for the feature tokens reduces $G$ to $Z^{\prime}=256$; the second linear layer reduces $V$ to $N$ then reshapes the tensor to ($B$, $N$, $F^{\prime\prime}=16$, $Z^{\prime\prime}=16$); the third linear layer changes $Z^{\prime\prime}$ to $Z=256$, then the last linear layer changes $F^{\prime\prime}$ to $F^{\prime}=128$. Thus, a tensor with shape ($B$, $N$, $F^{\prime}$, $Z$) is obtained. When using both the top and middle layers of the SoniDo features, we form such tensors for each layer following the pipeline above. The tensor(s) and output of the first encoder are then concatenated on the third axis. Finally, the size of the concatenated tensor is reduced to $256$ ($F$ in (Toyama et al., 2023)) by a linear layer. These $F^{\prime}$, $F^{\prime\prime}$, $Z^{\prime}$, and $Z^{\prime\prime}$ were determined from preliminary experiments.
We trained the following models that have different inputs: (1) spectrogram only, (2) spectrogram and the top layer of the SoniDo features, (3) spectrogram and the middle layer of the SoniDo features, (4) spectrogram and both the top and middle layer of tthe SoniDo features, (5) spectrogram and the MusicGen Large features, and (6) spectrogram and the MusicGen Small features, the same as the experiment described in D.1. We trained the models for 50 epochs on one A100 GPU. For the other conditions, we followed (Toyama et al., 2023). To confirm if the SoniDo and MusicGen features are useful when there are less training data, we train the models using 100, 50, 25, and 10% of training data. We chose the checkpoint with the highest F1 score in the validation split for further evaluation.
Figure 7: Modified version of hFT-Transformer for SoniDo and MusicGen feature tokens
@@
Figure 8 shows the loss curves of training and validation. The models using the SoniDo or MusicGen features reached a lower loss value at an earlier epoch compared with the baseline model using the spectrogram only as input. Table 13 lists the scores on the test set of MAPS. When all the training data were available, the models using the SoniDo and MusicGen features, except the model using the MusicGen Large features outperformed the baseline. When the models were trained with scarce data, the performance of the models using the SoniDo and MusicGen features was superior to that of the model using the spectrogram only. When the training data size was 50 and 25%, the performance of the models using the SoniDo features was still comparable to the baseline model trained with 100% data.
@@ E.3 Experiments: HTDemucs with SoniDo
  • HTDemucs (default): Model with default settings (Dora***https://github.com/facebookresearch/dora signature ‘955717e8’). The default batch size is $32$ which corresponds to $4$ samples per GPU as we trained in parallel on $8$ GPUs. The default number of training epochs is $360$.
@@
  • HTDemucs + CLAP: Same training settings as HTDemucs + SoniDo but where we used the embeddings from CLAP (Wu* et al., 2023) instead of the SoniDo features. More specifically, we used the ‘fine-grained embeddings‘ before the final AP in CLAP. This experiment enabled us to compare features obtained with two trained models (as opposed to the STFT)***Note that SoniDo and CLAP were trained on different datasets..

### v3

@@ 4.2.3 Music Mixing: Mix-Wave-U-Net, CRAFx2
Mix-Wave-U-Net (Steinmetz et al., 2022) along with a modified CRAFx (Martínez-Ramírez et al., 2020), henceforth referred to as CRAFx2, are used as the baselines. The input to both networks is the stereo stems pre-processed by Fx-normalization (Martínez-Ramírez et al., 2022), and the output is the stereo mixture. These models do not handle high-level information relevant to mixing, such as genre, instrumentation, or mood. Conditioning these models with extracted features, which implicitly contain such information, is expected to improve mixing performance. The features are computed from the monaural downmix of the mixture, which corresponds to the summation of the Fx-normalized input stems. To incorporate these features, we condition both networks using Feature-wise Linear Modulation (FiLM) layers (Perez et al., 2018). For Mix-Wave-U-Net, we inject features into the up-sampling and bottleneck one-dimensional (1D) convolutional blocks. For CRAFx2, we use FiLM layers to condition both the latent-space mixer and synthesis back-end (see Appendix F.1).
We train all models on MUSDB18, and the stereo-invariant loss, along with all training hyperparameters, remains the same, as in a previous study (Martínez-Ramírez et al., 2022). Due to the inherent subjectivity of the task, identifying the best model is challenging. Thus, as shown in Table 1, an objective evaluation is conducted by measuring the proximity between the output mixes and target mixes on the same test sets as in Martínez-Ramírez et al. (2022). The proximity measurement is based on objective metrics (Steinmetz et al., 2022; Martínez-Ramírez et al., 2022). These metrics consist of spectral, panning, dynamic, and loudness low-level audio features, which are the key audio characteristics often manipulated during the mixing process. Further details and experiments are provided in Appendices F.2 and F.3, respectively.
@@
Using the SoniDo features leads to the best performance. The Jukebox and MusicGen features provide improvements but not as effective as those of SoniDo, particularly for CRAFx2, where the default model outperforms both Jukebox and MusicGen in various metrics. For MusicGen, we assume this performance gap may be attributed to its training on 32 kHz audio compared with the 44.1 kHz used for SoniDo, which limit its effectiveness for full-band tasks, such as music mixing, that require higher sampling rates. There is no data-driven approach that used task-agnostic features of the input stems for music mixing improvement. Thus, we can conclude that incorporating the SoniDo features benefits both the training and performance of automatic music mixing models. This aligns with recent design studies (Lefford et al., 2021; Vanka et al., 2023), advocating for the incorporation of contextual inputs.
Table 6: Music mixing. Evaluation results on the MDXDB21-dry and MUSDB18 test sets include mean absolute percentage error for audio effect-related features, their average, and stereo-invariant loss. More details are provided in Table 15.
@@
Mix-Wave-U-Net (default) 0.234 0.215 0.073 0.168 0.173 89.631 0.201 0.164 0.085 0.167 0.154 34.253
Mix-Wave-U-Net + Jukebox 0.240 0.231 0.075 0.154 0.175 83.717 0.206 0.187 0.082 0.157 0.158 32.573
@@
Mix-Wave-U-Net + SoniDo 0.226 0.180 0.067 0.131 0.151 78.180 0.186 0.175 0.063 0.179 0.151 30.116
CRAFx2 (default) 0.193 0.179 0.070 0.152 0.148 82.095 0.193 0.154 0.081 0.165 0.148 32.856
CRAFx2 + Jukebox 0.231 0.249 0.075 0.144 0.175 87.973 0.216 0.220 0.082 0.165 0.171 36.172
@@ 6 Conclusions
We extended the use of music foundation models from MIR to generic music downstream tasks. The task-agnostic intermediate representation extracted using proposed SoniDo model has been applied to task-specific models of music tagging, music transcription, music source separation, and music mixing. On the basis of the evaluation results, performance improvement is observed for all selected music downstream tasks. This suggests that incorporating the intermediate features extracted from a pre-trained auto-regressive music foundation model should be considered as a generic booster in future development of task-specific models. This is especially helpful when it is difficult to acquire a sufficient dataset or the computation resource does not allow large-scale training. A study on the bias propagation of a pre-trained music foundation model to a downstream task model should be conducted in another future work.
Acknowledgement
We would like to thank Takashi Shibuya, Naoya Takahashi, Marc Ferras, and Masato Ishii for many helpful comments during the preparation of this manuscript. Besides, we thank anonymous reviewers for their valuable suggestions and comments.
@@ Appendix B Architecture of SoniDo
SoniDo is a generative model that generates music samples conditioned on given text prompts. We train the model using an internal dataset contains around 115k studio-quality library music tracks sampled at 44.1 kHz. Their lengths vary from 30s to 150s, their tempos vary between 50bpm to 200bpm, and the total length is around 4,000h. 90% of the dataset is non-vocal. Although there are more than 50 genres included in the dataset, it is biased toward orchestral and western music. To generate music samples, we use the three priors in SoniDo consecutively from the top level to the bottom level. It is important to note that during music generation, we use the token sequences sampled from top and middle priors for conditioning. For training and feature extraction, however, we rely on ground-truth tokens.
@@ B.3 Evaluation on Music Generation
For evaluation, we used the MusicCaps dataset (Agostinelli et al., 2023), which includes 5,521  [^2] We used a subset of 5,514 samples due to the presence of expired or invalid URLs in MusicCaps items. Efforts were made to retrieve the missing files; however, three samples are absent despite our efforts. pairs of audio clips and their text captions written by musicians. Subsequently, we used SoniDo to generate $10$-s audio samples, conditioned upon textual captions of MusicCaps. Following the objective evaluation methodology(Copet et al., 2023), we evaluated SoniDo’s performance on MusicCaps using three objective metrics: the Fréchet audio distance (FAD) (Kilgour et al., 2019), Kullback-Leibler Divergence (KL), and CLAP (Wu* et al., 2023) score.
@@
We used FAD to measure the overall quality of the generated samples. It is a reference-free metric that compares embedding statistics computed on an evaluation set with embedding statistics computed on the clean dataset. A lower FAD indicates a higher degree of similarity between the embedding statistics from generated audio samples and the embedding statistics from the ground-truth (i.e., MusicCaps). We used the official VGGish based-FAD computation module [^3] https://github.com/google-research/google-research/tree/master/frechet_audio_distance to extract embeddings from audio. Since VGGish was trained using $16$-kHz audio, all the audio data were resampled to $16$ kHz.
@@
MusicGen-large (public  [^4] Publicly released MusicGen-large trained on non-vocal dataset: https://github.com/facebookresearch/audiocraft/blob/main/model_cards/MUSICGEN_MODEL_CARD.md, ) 32kHz 5.88 1.41 0.27
@@ D.1 Piano Transcription with Linear Layers
The mel spectrogram is a 3D tensor ($B$, $N$, $F$), where $B=8$ is the batch size, $N=256$ is the number of frames, and $F=256$ is the number of frequency bins. A linear layer converts the tensor into ($B$, $N$, $Z$) along the last axis, where $Z=512$ is a common size for latent embeddings. The SoniDo, MusicGen , and Jukebox features are formed as a 3D tensor ($B$, $V$, $G$), where $V$ is the number of tokens per $N$ frames (2,048 for the top layer of SoniDo and for Jukebox, 8,192 for the middle layer of SoniDo, and 299 for MusicGen), and $G$ is the embedding size of features (4,800 for the top layer of SoniDo and for Jukebox, 3,200 for the middle layer of SoniDo, 2,048 for MusicGen Large, and 1,024 for MusicGen Small). The first linear layer converts $G$ to $Z$, then the second linear layer converts $V$ to $N$ to make the tensor into the shape of ($B$, $N$, $Z$). The tensors are then concatenated along the last axis. If the top layer of SoniDo, middle layer of SoniDo, or MusicGen is used, the dimension of the last axis is $2Z$, whereas if both layers of SoniDo are used, the dimension is $3Z$. The following linear layer is used to convert $2Z$ or $3Z$ to $2P$, where $P=88$ is the number of pitches. Finally, the tensor is split into two ($B$, $N$, $P$) tensors followed by a sigmoid activation to indicate the estimated frame and onset information. We followed Kong et al. (2021) to extract the precise timing of onsets from datasets, in which we set the hyper-parameter to control the target sharpness $J$ to $3$. The loss function for frame and onset is binary cross entropy.
@@
(a) Linear transcription model using SoniDo or Jukebox
(b) hFT-Transformer using SoniDo or Jukebox
@@
We trained multiple models with the following input: (1) spectrogram only, (2) spectrogram and the top layer of the SoniDo features, (3) spectrogram and the middle layer of the SoniDo features, (4) spectrogram and both the top and middle layers of the SoniDo features, (5) spectrogram and the MusicGen Large features, (6) spectrogram and the MusicGen Small features, and (7) spectrogram and the Jukebox features. For each model, we trained for 50 epochs on one A100 graphics processing unit (GPU), using Adam (Kingma & Ba, 2015) optimizer with a learning rate of 1e-4. PyTorch ReduceLROnPlateu was used for learning-rate scheduling with default parameters. We chose to use the checkpoint with the highest F1 score in the validation split for each model.
As listed in Table 3, the model using the SoniDo features, the MusicGen features, or the Jukebox features outperformed the spectrogram-only baseline. In particular, the contribution of SoniDo’s middle layer was higher than that of the top layer. These results suggest that the SoniDo features are promising for music transcription tasks.
@@ D.2 Linear Instrument-agnostic Music Transcription
The detailed model architecture of hFT-Transformer has been described in (Toyama et al., 2023). Figures 6(b) and (d) show the feature-spectrogram alignment scheme for hFT-Transformer. The feature tokens and spectrogram cannot be perfectly aligned due to the different sampling frequencies in hFT-Transformer ($16$ kHz), SoniDo ($44.1$ kHz), and MusicGen ($32$ kHz). The $N$ frame in the spectrogram corresponds to roughly $707$ tokens of the top layer of the SoniDo features and of the Jukebox features, $,82$ 828 tokens of the middle layer of the SoniDo features, and $104$ tokens of the MusicGen features. Figure 7 shows the modified hFT-Transformer that accepts the SoniDo or MusicGen features as the additional input. Following the setup in D.1, the SoniDo and MusicGen feature tokens have three dimensions ($B$, $V$, $G$), where $B=8$ is the batch size, $V$ is the number of tokens per number of frames $N=128$, and $G$ is the embedding size of the feature tokens. The first linear layer for the feature tokens reduces $G$ to $Z^{\prime}=256$; the second linear layer reduces $V$ to $N$ then reshapes the tensor to ($B$, $N$, $F^{\prime\prime}=16$, $Z^{\prime\prime}=16$); the third linear layer changes $Z^{\prime\prime}$ to $Z=256$, then the last linear layer changes $F^{\prime\prime}$ to $F^{\prime}=128$. Thus, a tensor with shape ($B$, $N$, $F^{\prime}$, $Z$) is obtained. When using both the top and middle layers of the SoniDo features, we form such tensors for each layer following the pipeline above. The tensor(s) and output of the first encoder are then concatenated on the third axis. Finally, the size of the concatenated tensor is reduced to $256$ ($F$ in (Toyama et al., 2023)) by a linear layer. These $F^{\prime}$, $F^{\prime\prime}$, $Z^{\prime}$, and $Z^{\prime\prime}$ were determined from preliminary experiments.
We trained the following models that have different inputs: (1) spectrogram only, (2) spectrogram and the top layer of the SoniDo features, (3) spectrogram and the middle layer of the SoniDo features, (4) spectrogram and both the top and middle layer of tthe SoniDo features, (5) spectrogram and the MusicGen Large features, (6) spectrogram and the MusicGen Small features, and (7) spectrogram and the Jukebox features, the same as the experiment described in D.1. We trained the models for 50 epochs on one A100 GPU. For the other conditions, we followed (Toyama et al., 2023). To confirm if these features are useful when there are less training data, we train the models using 100, 50, 25, and 10% of training data. We chose the checkpoint with the highest F1 score in the validation split for further evaluation.
Figure 7: Modified version of hFT-Transformer for SoniDo, MusicGen , and Jukebox feature tokens
@@
Figure 8 shows the loss curves of training and validation. The models using the SoniDo, MusicGen or Jukebox features reached a lower loss value at an earlier epoch compared with the baseline model using the spectrogram only as input. Table 13 lists the scores on the test set of MAPS. When all the training data were available, the models using the SoniDo, MusicGen , and Jukebox features, except the model using the MusicGen Large features outperformed the baseline. When the models were trained with scarce data, the performance of the models using the SoniDo, MusicGen , and Jukebox features was superior to that of the model using the spectrogram only. When the training data size was 50 and 25%, the performance of the models using the SoniDo features was still comparable to the baseline model trained with 100% data.
@@ E.3 Experiments: HTDemucs with SoniDo
  • HTDemucs (default): Model with default settings (Dora [^5] https://github.com/facebookresearch/dora signature ‘955717e8’). The default batch size is $32$ which corresponds to $4$ samples per GPU as we trained in parallel on $8$ GPUs. The default number of training epochs is $360$.
@@
  • HTDemucs + CLAP: Same training settings as HTDemucs + SoniDo but where we used the embeddings from CLAP (Wu* et al., 2023) instead of the SoniDo features. More specifically, we used the ‘fine-grained embeddings‘ before the final AP in CLAP. This experiment enabled us to compare features obtained with two trained models (as opposed to the STFT) [^6] Note that SoniDo and CLAP were trained on different datasets..
[…]